# Mean-field dynamics of attractive resource interaction: From uniform to aggregated states

### v1

@@ 3.4 Criteria for optimizing limit coordinates
This implies that all parameters $c_{k}$ satisfy the condition $c_{k}>\Lambda$, $k=\overline{1,n}$.
Figure 5: Evolution of the vector coordinates from the initial state ${\bf p}^{0}=(0.072;0.183;0.141;0.115;0.030;0.030;0.011;0.1666;0.116;0.136)$ to the limit state ${\bf p}^{\infty}=(0.160;0.148;0.135;0.122;0.109;0.096;0.080;0.068;0.054;0.028)$ under the action of conflict transformation (4). All parameters $c_{k}$ satisfy the condition $c_{k}>\Lambda$, $k=\overline{1,10}$ The limit state is entirely determined by the set of constants $\{c_{i}\}_{i=1}^{10}$, illustrating the optimization principles from 810: the coordinate $p_{1}^{t}$ with the maximum $c_{1}=0.71$ converges to the highest value, while $p_{10}^{t}$ with the minimum $c_{10}=0.61$ converges to the lowest.
Proposition 8 (On maximizing a limiting coordinate).
For maximizing the limiting coordinate $p_{i}^{\infty}$ for some fixed index $i$, it is necessary and sufficient that:
The parameter $c_{i}$ is maximal among all parameters $c_{k}$: $c_{i}=\max\limits_{k\in\{1,\dots,n\}}\{c_{k}\}.$
@@
To maximize the limiting coordinate $p_{i}^{\infty}=1-\frac{\Lambda}{c_{i}}$, it is necessary to minimize the positive fraction $\frac{\Lambda}{c_{i}}$. This is equivalent to maximizing the denominator $c_{i}\sum\limits_{m=1}^{n}\frac{1}{c_{m}}$. Let’s expand this product:
@@
Proposition 9 (On minimizing a limiting coordinate).
To minimize the limiting coordinate $p_{j}^{\infty}$ (given $p_{j}^{\infty}>0$) for some fixed index $j$, it is necessary and sufficient that:
@@
To minimize $p_{j}^{\infty}=1-\frac{\Lambda}{c_{j}}$ (while keeping $p_{j}^{\infty}>0$), it is necessary for the fraction $\frac{\Lambda}{c_{j}}$ to be maximal, but less than 1. This is equivalent to minimizing the denominator $c_{j}\sum\limits_{m=1}^{n}\frac{1}{c_{m}}$, provided it remains strictly greater than n1n-1. The expression $1+c_{j}\sum\limits_{m\neq j}\frac{1}{c_{m}}$ will be minimal (given $p_{j}^{\infty}>0$) when $c_{j}$ is minimal (condition (a)), and the sum $\sum\limits_{m\neq j}\frac{1}{c_{m}}$ is minimal. The latter condition holds when the values $c_{m}$ (for $m\neq j$) are as large as possible (condition (b)). Thus, conditions (a) and (b) are necessary and sufficient for minimizing $p_{j}^{\infty}$ (given $p_{j}^{\infty}>0$). ∎
Proposition 10 (On simultaneous optimization of limiting coordinates).
@@
The parameter $c_{i}$ is maximal among all parameters $c_{k}$: $c_{i}=\max\limits_{k\in\{1,\dots,n\}}\{c_{k}\}$.
The parameter $c_{j}$ is minimal among all parameters $c_{k}$: $c_{j}=\min\limits_{k\in\{1,\dots,n\}}\{c_{k}\}$.
@@
Let $c_{i}=\max\limits_{k\in\{1,\dots,n\}}\{c_{k}\}$, $c_{j}=\min\limits_{k\in\{1,\dots,n\}}\{c_{k}\}.$ To minimize $\Lambda$ (which is beneficial for $p_{i}^{\infty}$), the sum $\sum_{m=1}^{n}\frac{1}{c_{m}}$ must be large. Since $c_{i}$ is large and $c_{j}$ is small, $1/c_{i}$ is small and $1/c_{j}$ is large. To maximize the sum, the terms $1/c_{k}$ for $k\notin\{i,j\}$ must also be large. This is achieved if $c_{k}$ for $k\notin\{i,j\}$ are small, i.e., close to $c_{j}$ (condition (c)).
@@
4 Future outlook: from fixed points to complex dynamics via time-delayed feedback
@@
A particularly well-studied and powerful mechanism for this is time-delayed feedback, famously introduced by Pyragas as a method for controlling chaos [36, 37]. While originally conceived for control, time-delayed feedback is also a potent source of complex dynamics. Introducing a delay $\tau$ transforms the system into an infinite-dimensional one, known to undergo bifurcations leading to stable limit cycles and chaos [38, 39, 40, 41, 42, 43, 44], providing a clear theoretical pathway from our fixed-point dynamics to the rich behaviors observed in real-world systems.
This dynamic can be modeled by introducing a time-delayed dependence on the component’s own share. The favorability coefficient $c_{i}$ at time $t$ is then defined as:
@@
b) Limit cycle (quasi-periodic), $\beta=1.5$
c) Limit cycle (periodic), $\beta=3$
Figure 6: Phase portrait (attractor in the $p_{1}^{t}$, $p_{2}^{t}$ space), where ${\bf p}^{0}=(0.25,0.26,0.24,0.25)$, $c=(0.9,0.85,0.95,0.8)$, $\tau=30$.
The introduction of this feedback mechanism reveals a rich spectrum of dynamical behaviors far beyond a single fixed point. The feedback strength $\beta$ acts as a critical bifurcation parameter, governing the system’s transition from fixed point dynamics to complex aperiodic behavior.
Numerical simulations demonstrate that as $\beta$ increases, the system undergoes a transition from stable fixed points (as in Fig. 6a) to stable periodic oscillations (limit cycles) and eventually to highly complex, aperiodic fluctuations (as illustrated in Fig. 6b, c and the bifurcation diagram in Fig. 7).
Figure 7: Bifurcation diagram of the system, $c=(0.9,0.85,0.95,0.8)$, $\tau=30$.
For moderate values of $\beta$, the system exhibits stable limit cycles. This regime represents a dynamic equilibrium where no single component achieves permanent dominance, leading instead to a perpetual and predictable periodic or quasi-periodic redistribution of resources (Fig. 6b, c).
However, when $\beta$ exceeds a critical threshold, the system’s behavior becomes aperiodic and unpredictable (Fig. 6b). The trajectory remains confined to a specific region of the phase space, which we provisionally describe as a complex aperiodic attractor. Its aperiodic nature suggests significant sensitivity to initial conditions.
@@ 5 Conclusion
This study provides a comprehensive analysis of a dynamical system that models the redistribution of a certain conserved quantity (e.g., resource, capital, or influence) in a discrete state space under the influence of attractive interaction. The core findings demonstrate that such systems converge to an equilibrium state, and the nature of this equilibrium-uniform or aggregated-is critically dependent on the interaction constants $c_{i}$.
When all parameters $c_{i}$, which can be interpreted as the favorability of conditions or the intrinsic weights of the components, are uniform, the system inevitably converges to a uniform distribution, which is the sole stable equilibrium point. This highlights a theoretical pathway to achieving equality in resource distribution within this model, which is important for the analysis of idealized social or economic systems.
The research is further extended by analyzing the influence of heterogeneous constants $c_{i}$ on the limit state, showing that these parameters directly determine the final distribution. The explicit form of the limit vectors reveals that higher values of $c_{i}$ correlate with a greater concentration of the resource in the corresponding component. Importantly, the study details the conditions under which a component’s share of the resource might diminish to zero, which can be interpreted as a complete displacement from the system or a loss of influence.
The proposed analytical criteria provide precise conditions for maximizing or minimizing specific coordinates in the limiting state, offering insights into how external conditions (represented by $c_{i}$) can be manipulated to achieve desired distribution outcomes. This offers valuable theoretical tools for managing systems exhibiting attractive interactions.
The developed mathematical framework, particularly the characterization of fixed points and their stability based on interaction parameters, offers a foundational contribution to interdisciplinary fields studying social and economic dynamics, providing tools for the analysis and prediction of the behavior of complex adaptive systems.

### v2

@@ 3.4 Criteria for optimizing extremal coordinates
This implies that all parameters $c_{k}$ satisfy the condition $c_{k}>\Lambda$, $k=1,\ldots,n$.
Figure 5: Evolution of the vector coordinates from the initial state ${\bf p}^{0}=(0.072;0.183;0.141;0.115;0.030;0.030;0.011;0.1666;0.116;0.136)$ to the equilibrium ${\bf p}^{\infty}=(0.160;0.148;0.135;0.122;0.109;0.096;0.080;0.068;0.054;0.028)$ under the action of conflict transformation (4). All parameters $c_{k}$ satisfy the condition $c_{k}>\Lambda$, $k=1,\ldots,10$. The equilibrium is entirely determined by the set of constants $\{c_{i}\}_{i=1}^{10}$, illustrating the optimization principles from 79: the coordinate $p_{1}^{t}$ with the maximum $c_{1}=0.71$ converges to the highest value, while $p_{10}^{t}$ with the minimum $c_{10}=0.61$ converges to the lowest.
Proposition 7 (On maximizing a coordinate of the equilibrium).
For maximizing the coordinate $p_{i}^{\infty}$ for some fixed index $i$, it is necessary and sufficient that:
The parameter $c_{i}$ is maximal among all parameters $c_{k}$: $c_{i}=\max\limits_{k\in\{1,\ldots,n\}}\{c_{k}\}.$
@@
To maximize the coordinate $p_{i}^{\infty}=1-\frac{\Lambda}{c_{i}}$, it is necessary to minimize the positive fraction $\frac{\Lambda}{c_{i}}$. This is equivalent to maximizing the denominator $c_{i}\sum\limits_{m=1}^{n}\frac{1}{c_{m}}$. Let’s expand this product:
@@
Proposition 8 (On minimizing a coordinate of the equilibrium).
To minimize the coordinate $p_{j}^{\infty}$ (given $p_{j}^{\infty}>0$) for some fixed index $j$, it is necessary and sufficient that:
@@
To minimize the coordinate $p_{j}^{\infty}=1-\frac{\Lambda}{c_{j}}$ (while keeping $p_{j}^{\infty}>0$), it is necessary for the fraction $\frac{\Lambda}{c_{j}}$ to be maximal, but less than 1. This is equivalent to minimizing the denominator $c_{j}\sum\limits_{m=1}^{n}\frac{1}{c_{m}}$, provided it remains strictly greater than n1n-1. The expression $1+c_{j}\sum\limits_{m\neq j}\frac{1}{c_{m}}$ will be minimal (given $p_{j}^{\infty}>0$) when $c_{j}$ is minimal (condition (a)), and the sum $\sum\limits_{m\neq j}\frac{1}{c_{m}}$ is minimal. The latter condition holds when the values $c_{m}$ (for $m\neq j$) are as large as possible (condition (b)). Thus, conditions (a) and (b) are necessary and sufficient for minimizing $p_{j}^{\infty}$ (given $p_{j}^{\infty}>0$). ∎
Proposition 9 (On simultaneously optimizing the coordinates of the equilibrium).
@@
The parameter $c_{i}$ is maximal among all parameters $c_{k}$: $c_{i}=\max\limits_{k\in\{1,\ldots,n\}}\{c_{k}\}$.
The parameter $c_{j}$ is minimal among all parameters $c_{k}$: $c_{j}=\min\limits_{k\in\{1,\ldots,n\}}\{c_{k}\}$.
@@
Let $c_{i}=\max\limits_{k\in\{1,\ldots,n\}}\{c_{k}\}$, $c_{j}=\min\limits_{k\in\{1,\ldots,n\}}\{c_{k}\}.$ To minimize $\Lambda$ (which is beneficial for $p_{i}^{\infty}$), the sum $\sum_{m=1}^{n}\frac{1}{c_{m}}$ must be large. Since $c_{i}$ is large and $c_{j}$ is small, $1/c_{i}$ is small and $1/c_{j}$ is large. To maximize the sum, the terms $1/c_{k}$ for $k\notin\{i,j\}$ must also be large. This is achieved if $c_{k}$ for $k\notin\{i,j\}$ are small, i.e., close to $c_{j}$ (condition (c)).
@@
4 Future outlook: from fixed points to complex dynamics via delayed feedback
@@
A particularly well-studied and powerful mechanism for this is delayed feedback, famously introduced by Pyragas as a method for controlling chaos [40, 41]. While originally conceived for control, delayed feedback is also a potent source of complex dynamics. Introducing a delay $\tau$ transforms the system into an infinite-dimensional one, known to undergo bifurcations leading to stable limit cycles and chaos [42, 43, 44, 51, 52, 53]. Such complex mean–field dynamics provide a clear theoretical pathway from our fixed-point dynamics to the rich behaviors observed in macroscopic real–world systems, including biophysically grounded neural populations [54, 55] and complex quantum states [56].
This dynamic can be modeled by introducing a delayed dependence on the component’s own share. The favorability coefficient $c_{i}$ at time $t$ is then defined as:
@@
b) Quasi-periodic orbit, $\beta=1.5$
c) Periodic orbit, $\beta=3$
Figure 6: Phase portrait (attractor in the $p_{1}^{t}$, $p_{2}^{t}$ space), where ${\bf p}^{0}=(0.25,0.26,0.24,0.25)$, ${\bf c}=(0.9,0.85,0.95,0.8)$, $\tau=30$.
The introduction of this feedback mechanism reveals a rich spectrum of dynamical behaviors far beyond a single fixed point. Numerical simulations demonstrate that the feedback strength $\beta$ acts as a critical bifurcation parameter, driving the system’s complexity and governing the transition from fixed-point dynamics to complex aperiodic behavior. The bifurcation diagram (Fig. 7) is constructed by plotting the asymptotic values of the trajectory’s first coordinate, $p_{1}^{t}$, after discarding initial transient dynamics.
For moderate values of $\beta$, the system exhibits periodic and quasi-periodic oscillations (Fig. 6b, c). In terms of the resource competition model, this mathematical regime represents a dynamic equilibrium where no single component achieves permanent dominance, leading instead to a perpetual and predictable periodic or quasi-periodic redistribution of resources.
As $\beta$ increases past approximately 1.5, we phenomenologically observe that the system loses its stable fixed point and transitions into a broad regime of highly complex, aperiodic, and unpredictable oscillations. The trajectory remains confined to a specific region of the phase space, which we provisionally describe as a complex aperiodic (chaotic-like) attractor, suggesting significant sensitivity to initial conditions. These dynamics remain highly complex throughout a wide parametric window $\beta\in[1.5,4.1]$, where simulations reveal a dense bifurcation structure interspersed with sudden topological changes.
Figure 7: Bifurcation diagram of the extended system with delayed feedback for ${\bf c}=(0.9,0.85,0.95,0.8)$ and $\tau=30$. Asymptotic values of $p_{1}^{t}$ (after discarding $10^{4}$ transient steps) are plotted as a function of $\beta$, revealing a stable fixed point for $\beta\lessapprox 1.5$, a broad complex regime for $\beta\in[1.5,4.1]$, and a return to a stable fixed point following a global crisis near $\beta\approx 4.1$.
While a rigorous topological classification of these bifurcations falls outside the scope of this paper, these numerical results clearly confirm that delayed attractive interaction can drive the considered system through highly unpredictable oscillatory regimes and abrupt structural shifts, providing a rich ground for future analytical studies.
@@ 5 Conclusion
This study provides a comprehensive analysis of a dynamical system that models the redistribution of a conserved quantity (e.g., resource, capital, or influence) in a discrete state space under the influence of attractive interaction. The core mathematical findings demonstrate that trajectories of this system always converge to a stable fixed point. The topological location of this equilibrium — whether it lies in the interior of the simplex (all $p_{i}^{\infty}>0$) or on its boundary (some $p_{i}^{\infty}=0$) — is analytically determined by the interaction constants $c_{i}$. In terms of the applied model, this dictates whether the system reaches a state of coexistence or heavily aggregated resource concentration.
When all parameters $c_{i}$ are equal, all trajectories with strictly positive initial coordinates converge to a unique attractor located exactly at the barycenter of the simplex ($p_{i}^{\infty}=1/n$ for all $i$). This result admits a natural interpretation: it describes a theoretical pathway to achieving perfect equality in resource distribution, which is important for the analysis of idealized social or economic systems.
The research is further extended by analyzing the influence of heterogeneous constants $c_{i}$ on the equilibrium. The explicit analytical form of the fixed point reveals a strict functional dependence: higher values of a parameter $c_{i}$ strictly maximize the corresponding coordinate $p_{i}^{\infty}$. Furthermore, the study establishes exact bifurcation thresholds under which a specific coordinate asymptotically vanishes ($p_{i}^{\infty}=0$). In the context of resource competition, these mathematical phenomena represent, respectively, the systematic concentration of the resource and the complete competitive exclusion (displacement) of a component from the system.
The proposed analytical criteria translate these mathematical results into actionable conditions: by tuning the parameters $c_{i}$, one can deliberately steer the system toward a desired equilibrium configuration, whether maximizing a specific coordinate or driving others to zero. This offers a concrete theoretical tools for managing systems exhibiting attractive interactions.
The developed framework — comprising explicit equilibrium formulas, bifurcation thresholds, and stability criteria — provides a self–contained mathematical foundation for the analysis of discrete–time attractive interaction models, with direct applications to resource competition, opinion dynamics, and wealth distribution.